Secret:

Write the quadratic equation in standard form. (x - 3)^2 = 0 x^2 + 9 = 0 x^2 + 6x + 9 = 0 x^2 - 6x + 9 = 0

Hero:

@Secret you were given \((x - 3)^2 = 0\) and asked to write the equation in standard form. Any ideas on how we might get started with doing so?

Flamo:

Multiply (x - 3) by Its self? So It would be x^2 + 9 = 0?

Secret:

Ye?

Hero:

Actually, that is not correct. The proper way to get started with this is to multiply the binomials. Understand that if you are given a binomial square such as \((x - a)^2\), that expression can be expanded to \((x - a)(x - a)\). From there you can use the distributive property to expand the expression then simply it until you have a quadratic expression.

Hero:

simplify*

Secret:

Ohhh

Flamo:

Oh, I see

Secret:

Now what?

Hero:

You re-write the equation to reflect the above information. Can you try doing that here @Secret?

Secret:

(x - 3)^2 = 0 (x - 3)(x - 3) = 0 Like that??

Hero:

Yes correct. Now the next step would be to actually multiply the binomials.

Hero:

Would you like to try @Secret?

Secret:

(x - 3)(x - 3) = 0 x^2 - 3x -3x + 6 Like this?

Hero:

Question. How did you end up with 6 as the last term? Did you add or multiply to get that?

Secret:

i added....was i supposed to multiply?

Hero:

Yes.

Secret:

Oops...then it would be this: x^2 - 3x -3x + 9 Right?

Hero:

Yes, but the expression can be simplified further no?

Secret:

...idk

Hero:

Do you see any like terms in that expression?

Bob:

or add like terms or subtract them or what

Secret:

yes, there is -3x

Hero:

Okay, so that means -3x can be combined with -3x.

Secret:

So i should add them?

Hero:

Yes, correct. Go ahead and add them

Secret:

So it would be: x^2 - 6x + 9 Right?

Hero:

Correct.

Hero:

So which answer choice is correct?

Secret:

The last one! Thx

Hero:

You're most welcome. If you have any more questions, feel free to close this one and post a new question.

Secret:

^-^ Ok Thank you